# QoE Maximization for Multiple-UAV-Assisted Multi-Access Edge Computing: An Online Joint Optimization Approach

### v1

@@ 18      Update $l=l+1$;
5.2 Stage 2: SUAV Trajectory Planning
Given the optimal task offloading decisions $\mathcal{A}^{*}$ and resource allocation $\{\mathcal{F}^{*},\mathcal{W}^{*}\}$, while removing irrelevant constant terms, problem $\textbf{P}^{\prime}$ can be converted into subproblem $\mathbf{P2}$ to decide trajectory planning for SUAVs, which is expressed as follows:
$\mathbf{P2}:\underset{\mathcal{Q}}{\text{min}}\ \sum_{n\in\mathcal{N}}\sum_{m\in\mathbf{M}_{n}(\mathcal{A})}V\frac{\gamma_{m}^{\text{T}}D_{m}+\gamma_{m}^{\text{E}}p_{m}D_{m}}{w_{n,m}^{*}B_{n}\log_{2}(1+\frac{\phi_{n,m}}{(\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{m}\|^{2}+H^{2})})}+$
$\sum_{n\in\mathcal{N}}Q_{n}^{\text{p}}\left(C_{1}\left(1+\frac{3v_{n}^{2}}{U_{\text{p}}^{2}}\right)+C_{2}\sqrt{\sqrt{C_{3}+\frac{v_{n}^{4}}{4}}-\frac{v_{n}^{2}}{2}}+C_{4}v_{n}^{3}\right)\tau$ (52)
@@ 5.2 Stage 2: SUAV Trajectory Planning
where $\mathbf{q}_{n^{\prime}}=\mathbf{q}_{n}(t+1)$, $\mathbf{q}_{n}=\mathbf{q}_{n}(t)$, $v_{n}=\frac{\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{n}\|}{\tau}$, and $\phi_{n,m}=\frac{p_{m}10^{-\frac{20\log_{10}\left(\frac{4\pi f_{n}}{c}\right)+\rho_{n,m}^{\text{LoS}}\eta^{\text{LoS}}+\left(1-\rho_{n,m}^{\text{LoS}}\right)\eta^{\text{nLoS}}}{10}}}{\varpi_{0}}$. Clearly, the objective function in (52) is non-convex with respect to $\mathbf{q}_{n^{\prime}}$ due to the following non-convex terms
$\small\begin{cases}TMR_{n}=\sqrt{\sqrt{C_{3}+\frac{v_{n}^{4}}{4}}-\frac{v_{n}^{2}}{2}},\ \forall n\in\mathcal{N},\\ TML_{n,m}=\frac{1}{\log_{2}\left(1+\frac{\phi_{n,m}}{(\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{m}\|^{2}+H^{2})}\right)},\ \forall n\in\mathcal{N},m\in\mathbf{M}_{n}(\mathcal{A}).\end{cases}$ (53)
Therefore, it is difficult to directly solve problem P2. Next, we transform the objective function into a convex function by introducing slack variables.
@@
For the non-convex term $TML_{n,m}$, we introduce the slack variable $\zeta_{n,m}$ such that $\zeta_{n,m}=TML_{n,m}$ and add the following constraint:
$\small\zeta_{n,m}\leq\log_{2}\left(1+\frac{\phi_{n,m}}{\left(H^{2}+\left\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{m}\right\|^{2}\right)}\right),\forall n\in\mathcal{N},m\in\mathbf{M}_{n}(\mathcal{A}).$ (55)
@@
where $\mathbf{Q}=\{\mathbf{q}_{n^{\prime}}\}_{n\in\mathcal{N}}$, $\mathbf{\zeta}=\{\zeta_{n,m}\}_{n\in\mathcal{N},m\in\mathbf{M}_{n}(\mathcal{A})}$ and $\mathbf{\xi}=\{\xi_{n}\}_{n\in\mathcal{N}}$.
@@
Suppose $\{\mathcal{Q}^{*},\mathbf{\zeta}^{*},\mathbf{\xi}^{*}\}$ is the optimal solution of problem $\mathbf{P2}^{\prime}$. The following equation holds:
$\zeta_{n,m}^{*}=\log_{2}\left(1+\frac{\phi_{n,m}}{\left(H^{2}+\left\|\mathbf{q}^{*}_{u^{\prime}}-\mathbf{q}_{m}\right\|^{2}\right)}\right),\forall n\in\mathcal{N},m\in\mathbf{M}_{n},$ (57)
$\xi_{n}^{*}=\sqrt{\sqrt{C_{3}+\frac{(v^{*}_{n})^{4}}{4}}-\frac{(v^{*}_{n})^{2}}{2}},\forall n\in\mathcal{N},$
where $v^{*}_{n}=\frac{\|\mathbf{q}^{*}_{n^{\prime}}-\mathbf{q}_{n}\|}{\tau}$. Otherwise, we can further decrease the value of the objective function by choosing a smaller $\xi_{n}$ or a larger $\zeta_{n,m}$ without violating constraints (54) and (55). Therefore, $\mathcal{Q}^{*}$ is also the optimal solution to problem $\mathbf{P2}$.
For problem $\mathbf{P2}^{\prime}$, the optimization objective (56) is convex but constraints (54), (55) and (3c) are still non-convex. Similar to [43, 27], the successive convex approximation (SCA) method is adopted to solve the non-convexity of above constraints, as demonstrated in the following Theorems 6,  7, and 8.
@@
For constraint (54), let $f_{n}(\mathbf{q}_{n^{\prime}},\xi_{n})=\xi_{n}^{2}+v_{n}^{2}$ and given a local point $\mathbf{q}^{(l)}_{n^{\prime}}$ at the $l$-th iteration, we can obtain a global concave lower bound for $f_{n}(\mathbf{q}_{n^{\prime}},\xi_{n})$ as
@@
Since $f_{n}(\mathbf{q}_{n^{\prime}},\xi_{n})$ is a convex quadratic form, the first-order Taylor expansion of $f_{n}(\mathbf{q}_{n^{\prime}},\xi_{n})$ at local point $\mathbf{q}^{(l)}_{n^{\prime}}$ is a global concave lower bound.
@@
For constraint (55), let $g_{n,m}(\mathbf{q}_{n^{\prime}})=\log_{2}\left(1+\frac{\phi_{n,m}}{\left(H^{2}+\left\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{m}\right\|^{2}\right)}\right)$ and given a local point $\mathbf{q}^{(l)}_{n^{\prime}}$ at the $l$-th iteration, we can obtain a global concave lower bound for $g_{n,m}(\mathbf{q}_{n^{\prime}})$ as
$g_{n,m}^{(l)}(\mathbf{q}_{n^{\prime}})\triangleq\log_{2}\left(1+\frac{\phi_{n,m}}{\left(H^{2}+\|\mathbf{q}_{n^{\prime}}^{(l)}-\mathbf{q}_{m}\|^{2}\right)}\right)$
@@
We first consider the function $f(x)=\log_{2}(1+\frac{\phi}{(H^{2}+x)})$, where $\phi>0$, and $x\geq 0$. The first-derivative of $f(x)$ can be calculated as:
$\frac{\partial f(x)}{\partial x}=-\frac{\phi(\log_{2}e)}{[\phi+(H^{2}+x)](H^{2}+x)}.$ (61)
Accordingly, the second-order derivative of $f(x)$ can be calculated as:
$\frac{\partial^{2}f(x)}{\partial x^{2}}=\frac{\phi(\log_{2}e)[2(H^{2}+x)+\phi]}{[\phi(H^{2}+x)+(H^{2}+x)^{2}]^{2}}.$ (62)
Clearly, since $\frac{\partial^{2}f(x)}{\partial x^{2}}>0$, $f(x)$ is a convex function. Therefore, the first-order Taylor expansion of $f(x)$ at a local point $x_{0}$ is a global concave under-estimator of $f(x)$, i.e., the following inequality holds for any $x$
$\small f(x)\geq\log_{2}(1+\frac{\phi}{(H^{2}+x_{0})})-\frac{\phi(\log_{2}e)(x-x_{0})}{[\phi+(H^{2}+x_{0})](H^{2}+x_{0})}.$ (63)
Substituting $\phi=\phi_{n,m}$, $x=\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{m}\|^{2}$ and $x_{0}=\|\mathbf{q}^{(l)}_{n^{\prime}}-\mathbf{q}_{m}\|^{2}$ into the inequality (63), we can prove the theorem.
@@
For constraint (3c), let $h_{i,j}(\mathbf{q}_{i^{\prime}},\mathbf{q}_{j^{\prime}})=\|\mathbf{q}_{i^{\prime}}-\mathbf{q}_{j^{\prime}}\|^{2}$ and given a local point $(\mathbf{q}_{i^{\prime}}^{(l)},\mathbf{q}_{j^{\prime}}^{(l)})$ at the $l$-th iteration, we can obtain a global concave lower bound for $h_{i,j}(\mathbf{q}_{i^{\prime}},\mathbf{q}_{j^{\prime}})$ as
@@
Since the Hessian matrix of $h_{i,j}(\mathbf{q}_{i^{\prime}},\mathbf{q}_{j^{\prime}})$ is positive semidefinite, $h_{i,j}(\mathbf{q}_{i^{\prime}},\mathbf{q}_{j^{\prime}})$ is a convex function. Therefore, the first-order Taylor expansion of $h_{i,j}(\mathbf{q}_{i^{\prime}},\mathbf{q}_{j^{\prime}})$ at local point $(\mathbf{q}_{i^{\prime}}^{(l)},\mathbf{q}_{j^{\prime}}^{(l)})$ is a global concave lower bound.
According to Theorems 6, 7, and 8, at the $l$-th iteration, constraints (54), (55), and (3c) can be approximated as:
@@
which are convex. Therefore, problem $\mathbf{P2}^{\prime}$ is converted into a convex optimization problem, which can be efficiently resolved by off-the-shelf optimization tools such as CVX [52]. We summarize the second stage algorithm in Algorithm 2.
@@ 6       Update $l=l+1$;
5.3 Main Steps of OJTRTA and Performance Analysis
In this section, the main steps of OJTRTA are described in Algorithm 3, and the corresponding analyses are provided in Theorems 9, 10, and 11.
@@ 11       Update $t=t+1$;
According to Lemma 4.11 in [48], the T-slot drift-plus-penalty achieved by the proposed algorithm ensures that:
$\small L(\bm{\Theta}(T))-L(\bm{\Theta}(1))+V\sum_{t=1}^{T}C_{s}(t)\leq WT^{2}+CT+VTC_{s}^{\mathrm{opt}}.$ (69)
Using the fact that $L(\bm{\Theta}(T))\geq 0$ and $L(\bm{\Theta}(1))=0$, and dividing by $VT$ for the above inequality, we complete the proof of the theorem.
@@
According to Theorem 4.13 of [48], we can conclude that all virtual queues are rate stable. Therefore, we have
$\lim_{T\rightarrow+\infty}\frac{Q_{n}^{\text{c}}(T)+Q_{n}^{\text{p}}(T)}{T}=0\ \text{with probability 1},\ \forall n\in\mathcal{N}.$ (70)
Using the sample path property (Lemma 2.1 of  [48]), we have
$\small\frac{Q_{n}^{\text{c}}(T)+Q_{n}^{\text{p}}(T)}{T}-\frac{Q_{n}^{\text{c}}(1)+Q_{n}^{\text{p}}(1)}{T}\geq\frac{1}{T}\sum_{t=1}^{T}E_{n}(t)-\frac{1}{T}\sum_{t=1}^{T}\bar{E}_{n}.$ (71)
By taking the infinite limit on both sides of (71) and given that $Q_{n}^{\text{c}}(1)=0$, $Q_{n}^{\text{p}}(1)=0$, we can prove that formula (16) holds.
@@
The proposed OJTRTA has a polynomial worst-case complexity in each time slot, i.e., $\mathcal{O}(I_{c}MN+(N+M)^{2.5}(N^{2}+M)\log_{2}(\frac{1}{\varepsilon}))$, where $I_{c}$ represents the number of iterations required for Algorithm 1 to converge to the Nash equilibrium, $M$ denotes the number of UDs and $\varepsilon$ is the accuracy of SCA for solving problem $\mathbf{P2}^{\prime}$.
@@
OJTRTA contains two phases in each time slot, i.e., Algorithm 1 and Algorithm 2. In Algorithm 1, assuming that the outer iteration (i.e., Lines $2-16$) converges after $I_{c}$ iterations, the computational complexity can be calculated as $\mathcal{O}(I_{c}MN)$. In Algorithm 2, problem $\mathbf{P2}^{\prime}$ can contain at most $2N+M$ variables and $2N+M+\frac{N(N-1)}{2}$ constraints. According to the analysis in [30], the number of iterations required is $\sqrt{2N+M}\log_{2}(\frac{1}{\varepsilon})$, where $\varepsilon$ is the accuracy of SCA. Therefore, the computational complexity of Algorithm 2 can be calculated as $\mathcal{O}((2N+M)^{2}\sqrt{2N+M}\log_{2}(\frac{1}{\varepsilon})(2N+M+\frac{N(N-1)}{2}))$, which is equivalent to $\mathcal{O}((N+M)^{2.5}(N^{2}+M)\log_{2}(\frac{1}{\varepsilon}))$. As a result, the computational complexity of OJTRTA is $\mathcal{O}(I_{c}MN+(N+M)^{2.5}(N^{2}+M)\log_{2}(\frac{1}{\varepsilon}))$ in the worst case.
Accordingly, it is proven that the proposed approach can effectively guarantee the performance of the system and meet the SUAV energy consumption constraint with the feasible computational complexity.
6 Simulation Results

### v2

@@ 18      Update $l=l+1$;
(3) Analysis of Nash Equilibrium. Based on the above analysis, a Nash equilibrium can be obtained by the proposed algorithm. Next, we first analyze the computational complexity, which is presented in Theorem 5.
Theorem 5.
The number of iterations $I_{c}$ required for the proposed algorithm to converge to a Nash equilibrium is upper bounded as follows:
$I_{c}\leq\frac{M^{2}(\beta_{\text{max}}^{2}+\phi_{\text{max}}^{2})+M(Q_{\text{max}}+U_{\text{max}}^{\text{loc}})}{\epsilon_{\text{min}}}.$ (41)
Proof.
Refer to Section VI of the supplementary material.
To evaluate the performance of the equilibrium solution, the Price of Anarchy (PoA) is introduced to quantify the gap between the worst-case equilibrium and the centralized optimal solutions, which can provide a bound on the sub-optimality of our proposed algorithm. Mathematically, PoA can be defined as
$\mathrm{PoA}=\frac{\max_{\mathcal{A}\in\mathbf{A}^{*}}\sum_{m\in\mathcal{M}}U_{m}(\mathcal{A})}{\min_{\mathcal{A}\in\mathbb{A}}\sum_{m\in\mathcal{M}}U_{m}(\mathcal{A})},$ (42)
where $\mathbf{A}^{*}$ denotes the set of Nash equilibrium for the MUD-TOG. Clearly, a smaller PoA indicates better performance of the obtained equilibrium solution. In the following, we illustrate the bound of the PoA through Theorem 6.
Theorem 6.
Let $a_{m}=m$ indicate the local computing, $U(\mathcal{A})$ denote the objective function of problem P1.2, i.e., $U(\mathcal{A})=\sum_{m\in\mathcal{M}}U_{m}(\mathcal{A})$, $\mathcal{A}^{*}$ indicates a Nash equilibrium, and $\hat{\mathcal{A}}$ denotes the centralized optimal strategy. For the formulated MUD-TOG, the PoA defined in Eq. (42) satisfies
$1\leq\text{PoA}\leq 3-\frac{\sum_{i\in\mathcal{I}}\big{(}G_{i}(\mathcal{A}^{*})+G_{i}(\hat{\mathcal{A})}\big{)}}{U(\hat{\mathcal{A}})},$ (43)
where $\mathcal{I}=\mathcal{M}\cup\mathcal{N}\cup\{u\}$ represents the task offloading strategy space of all UDs, $G_{i}(\mathcal{A})=\sum_{m\in\mathbf{M}_{i}(\mathcal{A})}G_{i,m}$, and $\mathbf{M}_{i}(\mathcal{A})=\{m|a_{m}=i,\forall m\in\mathcal{M}\}$ denotes the set of UDs whose offloading decision satisfies $a_{m}=i$. Furthermore, $G_{i,m}=Q_{i}^{\text{c}}E_{i,m}^{\text{c}}/V$ if $i\in\mathcal{N}$; otherwise, $G_{i,m}=0$.
Proof.
Refer to Section VII of the supplementary material.
V-B Stage 2: SUAV Trajectory Control
Given the task offloading decisions $\mathcal{A}^{*}$ and resource allocation $\{\mathcal{F}^{*},\mathcal{W}^{*}\}$, while removing irrelevant constant terms, problem $\textbf{P}^{\prime}$ can be converted into subproblem $\mathbf{P2}$ to decide trajectory control for SUAVs, which is given as follows:
$\mathbf{P2}:\underset{\mathcal{Q}}{\text{min}}\ \sum_{n\in\mathcal{N}}\sum_{m\in\mathbf{M}_{n}(\mathcal{A})}V\frac{\gamma_{m}^{\text{T}}D_{m}+\gamma_{m}^{\text{E}}p_{m}D_{m}}{w_{n,m}^{*}B_{n}\log_{2}(1+\frac{\phi_{n,m}}{\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{m}\|^{2}+H^{2}})}+$
$\sum_{n\in\mathcal{N}}Q_{n}^{\text{p}}\left(C_{1}\big{(}1+\frac{3v_{n}^{2}}{U_{\text{p}}^{2}}\big{)}+C_{2}\sqrt{\sqrt{C_{3}+\frac{v_{n}^{4}}{4}}-\frac{v_{n}^{2}}{2}}+C_{4}v_{n}^{3}\right)\tau$ (44)
@@ V-B Stage 2: SUAV Trajectory Control
where $\mathbf{q}_{n^{\prime}}=\mathbf{q}_{n}(t+1)$, $\mathbf{q}_{n}=\mathbf{q}_{n}(t)$, $v_{n}=\frac{\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{n}\|}{\Delta t}$, and $\phi_{n,m}=\frac{p_{m}c^{2}[\rho_{n,m}^{\text{LoS}}|h_{n,m}^{\text{LoS}}|^{2}\eta^{\text{NLoS}}+(1-\rho_{n,m}^{\text{LoS}})|h_{n,m}^{\text{NLoS}}|^{2}\eta^{\text{LoS}}]}{(4\pi f_{n})^{2}\varpi_{0}\eta^{\text{LoS}}\eta^{\text{NLoS}}}$. Clearly, the objective function in (44) is non-convex with respect to $\mathbf{q}_{n^{\prime}}$ due to the following non-convex terms
$\small\begin{cases}TMR_{n}=\sqrt{\sqrt{C_{3}+\frac{v_{n}^{4}}{4}}-\frac{v_{n}^{2}}{2}},\ \forall n\in\mathcal{N},\\ TML_{n,m}=\frac{1}{\log_{2}\big{(}1+\frac{\phi_{n,m}}{\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{m}\|^{2}+H^{2}}\big{)}},\ \forall n\in\mathcal{N},m\in\mathbf{M}_{n}(\mathcal{A}).\end{cases}$ (45)
Therefore, solving problem P2 directly is difficult. Next, we transform the objective function into a convex function by introducing slack variables.
@@
For the non-convex term $TML_{n,m}$, we introduce the slack variable $\zeta_{n,m}$ such that $1/\zeta_{n,m}=TML_{n,m}$ and add the following constraint:
$\small\zeta_{n,m}\leq\log_{2}\big{(}1+\frac{\phi_{n,m}}{H^{2}+\left\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{m}\right\|^{2}}\big{)},\forall n\in\mathcal{N},m\in\mathbf{M}_{n}(\mathcal{A}).$ (47)
@@
where $\mathcal{Q}=\{\mathbf{q}_{n^{\prime}}\}_{n\in\mathcal{N}}$, $\mathbf{\zeta}=\{\zeta_{n,m}\}_{n\in\mathcal{N},m\in\mathbf{M}_{n}(\mathcal{A})}$ and $\mathbf{\xi}=\{\xi_{n}\}_{n\in\mathcal{N}}$.
@@
Refer to Section VIII of the supplementary material.
For problem $\mathbf{P2}^{\prime}$, the optimization objective (48) is convex but constraints (46), (47) and (3c) are still non-convex. Similar to [44, 28], the successive convex approximation (SCA) method is adopted to solve the non-convexity of above constraints, as demonstrated in the Theorems 8,  9, and 10.
@@
For constraint (46), let $f_{n}(\mathbf{q}_{n^{\prime}},\xi_{n})=\xi_{n}^{2}+v_{n}^{2}$ and given a local point $\mathbf{q}^{(l)}_{n^{\prime}}$ at the $l$-th iteration, a global concave lower bound of $f_{n}(\mathbf{q}_{n^{\prime}},\xi_{n})$ can be obtained as
@@
Refer to Section IX of the supplementary material.
@@
For constraint (47), let $g_{n,m}(\mathbf{q}_{n^{\prime}})=\log_{2}\big{(}1+\frac{\phi_{n,m}}{H^{2}+\left\|\mathbf{q}_{n^{\prime}}-\mathbf{q}_{m}\right\|^{2}}\big{)}$ and given a local point $\mathbf{q}^{(l)}_{n^{\prime}}$ at the $l$-th iteration, a global concave lower bound of $g_{n,m}(\mathbf{q}_{n^{\prime}})$ can be obtained as follows:
$g_{n,m}^{(l)}(\mathbf{q}_{n^{\prime}})\triangleq\log_{2}\big{(}1+\frac{\phi_{n,m}}{H^{2}+\|\mathbf{q}_{n^{\prime}}^{(l)}-\mathbf{q}_{m}\|^{2}}\big{)}$
@@
Refer to Section X of the supplementary material.
@@
For constraint (3c), let $h_{i,j}(\mathbf{q}_{i^{\prime}},\mathbf{q}_{j^{\prime}})=\|\mathbf{q}_{i^{\prime}}-\mathbf{q}_{j^{\prime}}\|^{2}$ and given a local point $(\mathbf{q}_{i^{\prime}}^{(l)},\mathbf{q}_{j^{\prime}}^{(l)})$ at the $l$-th iteration, a global concave lower bound of $h_{i,j}(\mathbf{q}_{i^{\prime}},\mathbf{q}_{j^{\prime}})$ can be obtained as follows:
@@
Refer to Section XI of the supplementary material.
According to Theorems 8, 9, and 10, at the $l$-th iteration, constraints (46), (47), and (3c) can be approximated as
@@
which are convex. Therefore, problem $\mathbf{P2}^{\prime}$ is converted into a convex optimization problem, which can be efficiently resolved by off-the-shelf optimization tools like CVX [51]. We summarize the second stage algorithm in Algorithm 2.
@@ 6       Update $l=l+1$;
V-C Main Steps of OJTRTA and Performance Analysis
In this section, the main steps of OJTRTA are described in Algorithm 3, and the corresponding analyses are provided in Theorems 11, 12, and 13.
Specifically, at the beginning of each time slot, UDs first upload their task request information (line 3). Based on the collected UD information, the proposed approach derives the strategies of resource allocation and task offloading by invoking Algorithm 1 (line 4). Given these strategies, the proposed approach then determines the trajectory control for the SUAVs by invoking Algorithm 2 (line 5). Subsequently, the UDs execute their computational tasks according to the established offloading strategies, while the SUAVs allocate computing and communication resources to provide edge computing services and move to the next location (lines 6 and 7). Finally, the system cost and energy consumption queues are calculated and updated (lines 8 to 10).
@@ 11       Update $t=t+1$;
Refer to Section XII of the supplementary material.
@@
Refer to Section XIII of the supplementary material.
@@
The proposed OJTRTA has a polynomial worst-case complexity in each time slot, i.e., $\mathcal{O}(I_{c}MN+(N+M)^{2.5}(N^{2}+M)\log_{2}(\frac{1}{\varepsilon}))$, where $M$ represents the number of UDs, $I_{c}$ denotes the number of iterations necessary for Algorithm 1 to reach convergence at the Nash equilibrium, and $\varepsilon$ is the accuracy of SCA for solving problem $\mathbf{P2}^{\prime}$.
@@
Refer to Section XIV of the supplementary material.
Note that the proposed approach strikes a trade-off between computational complexity and performance to ensure feasibility, which is reasonable. Specifically, for the complex JTRTOP, obtaining the optimal solution usually incurs a large computational overhead or is even impossible, making it unsuitable for the real-time decision-making scenario considered. In contrast, the proposed approach reduces computational complexity at the expense of some performance trade-offs, which is realistic since the complexity reduction often comes with performance degradation [52]. Despite these trade-offs, the proposed approach enables real-time decision-making and achieves a feasible solution that satisfies the requirements of UDs while meeting the constraints of the system. Moreover, the simulation results will demonstrate the superior performance of the proposed approach.
VI Simulation Results